# SmartPlay : A Benchmark for LLMs as Intelligent Agents

### v1

@@ 1 Introduction
We identify $4$ key challenges important for general intelligent LLM agents but not captured in previous benchmarks. First, lots of real-world tasks require an agent to do long-horizon planning and execution. Second, many events are probabilistic and an intelligent agent may be expected to understand the odds. Third, we live in a 3D world, and many real-world tasks require an agent’s spatial reasoning. Fourth, when encountered with unseen situations, an intelligent agent should be able to learn from interactions or mistakes.
@@ 2.1 Research Challenges
The SmartPlay benchmark encapsulates a diverse set of challenges that evaluate various AI capabilities, as itemized in Table 2. For instance, Bandits primarily focuses on simple randomness, requiring minimum text understanding and rule-following. On the other hand, Rock Paper Scissors uniquely puts an emphasis on randomness and multiple game rules. Hanoi presents an advanced setting for object dependency reasoning, strategic planning, and handling mistakes. Messenger puts challenge on 2D spatial reasoning, reading syntactic variations and conducting multi-hop reasoning. Meanwhile, Minecraft offers a unique challenge in 3D spatial reasoning and generalization within a randomized world. We hope the SmartPlay benchmark would serve as a tool for identifying these nuanced gaps and directing future research.
@@
Reasoning with object dependencies 1. $0\sim 1$-hop dependency 2. $2\sim 3$-hop dependencies 3. multi-hop dependencies
Instruction/Rule Following 1. single game rule 2. <3<3 game rules 3. $5+$ game rules
Planning Ahead 1. <5<5 planning steps 2. $5+$ planning steps 3. concurrent objectives
@@
Randomness 1. no randomness 2. randomness present in game 3. randomness as core mechanism
@@
Error/Mistake Handling 1. not required 2. rollback 3. reasoning and learning
@@
Figure 2: We identify a set of 9 important capabilities for an intelligent agent. We identify different degrees of challenge for each capability as shown on the left. Each game in SmartPlay challenges unique set of capabilities at different degrees, as shown in the spider charts. We numerical values fo the spider plots in Table 3.
@@ 2.2 Two Armed Bandits
The LLM agent is provided two slot machines with hidden pre-defined reward probabilities $p_{1},p_{2}$. For slot machine $i$, the reward for the two possible out-comes are: $r_{i}$ for pay-off event and $r_{i}$ for no-pay-off event. The goal of the game is to find the arm with better return and maximize the reward over the course of 50 rounds.
@@
The human written manual informs the LLM of the number of slot machines (two) and the game objective.
@@ 2.3 Rock Paper Scissors
This benchmark follows the exact same game rules as the famous zero-sum game Rock Paper Scissors222wikipedia.org/wiki/Rock_paper_scissors.
The LLM agent plays against a hand-coded opponent that follows a hidden pre-defined strategy with probabilities $p_{1},p_{2},p_{3}$ for rock, paper, and scissors respectively. The scores for winning under each action is pre-defined and revealed to the LLM as $s_{1},s_{2},s_{3}$.
An agent must keep track of win/losses from its past roll-outs to analyze the opponent behavior, and then exploit the opponent to maximize payoff. Overall, the challenges include: 1) long context understanding, 2) understanding randomness, 3) learning from interactions, and 4) instruction following and rule parsing.
The human written manual provides instruction on the possible actions and how the win/draw/lose of each round is calculated.
@@ 2.4 Tower of Hanoi
The Tower of Hanoi requires the agent to think strategically and plan ahead, and put strict requirements on the LLM agents’ ability to understand and follow the rules of the game. The game can become more challenging if the agent makes a mistake. Sometimes, an agent may have to undo several moves to correct an error. Overall, the challenges include: 1) planing, 2) reasoning with dependencies, 3) recovery from mistakes.
@@ 2.5 MESSENGER
To succeed in MESSENGER, an agent must first relate entities and dynamics of the environment to their reference synonyms in the manual, identify message and goal objects, and navigate to bring the message to the goal while avoiding the enemy. The manual, by design, is challenging to understand even for human readers. Level 1 primarily challenges the agent’s 1) natural language understanding and 2) generalization . Level 2 includes additional challenge on the agent’s 3) reasoning with dependencies and 4) 2D spatial reasoning. Level 3 increases difficulty by adding distraction objects.
The original manuals provided by Hanjie et al. (2021) contain descriptions of the entities and world dynamics obtained through crowd-sourced human writers. We augment the manual with a specification on the game objective, and an “advice” for LLM agent to first identify goal objects and then approach its objective. The “advice” helps to guide LLM in approaching the particularly game manual.
@@ 2.6 Crafter
To succeed in Crafter an LLM agent has to first understand and master a variety of reusable skills composed of 17 actions. The agent needs to learn to navigate through up to 5 2D terrains (biomes), avoiding obstacles and dangerous creatures. The agent also needs to collect different resources and craft more advanced weapons/tools to unlock more skills and achievements, while at the same time balance crafting goals with survival goals like maintaining health, thirst, food, and rest (Hafner, 2021). Overall, the challenges include: 1) 2D spatial reasoning, 2) Changing environment dynamics, 3) Long context understanding, 4) Planning ahead, 5) Generalization, 6) Correcting from mistakes.
We provide the “context” string from Wu et al. (2023c) as the manual, generated by parsing the source-code of (Hafner, 2021). The “context” string has been shown to greatly improve performance of GPT-4 and text-davinci-003 on Crafter (Wu et al., 2023c). Interestingly, the “context” string does not capture all information necessary to succeed in the game, i.e., it requires 2 woods to craft the crafting table, and 8 stones to craft the furnance. The agent has to 7) learn from interaction.
@@ 2.7 Minecraft
To succeed in the creative “find” tasks, a LLM agent has to have enough domain knowledge about different biomes in Minecraft, and be able to correlate visual observation (text description of visual world) with domain knowledge, and navigate in a 3D environment. Overall, the challenges include: 1) planning ahead, 2) domain knowledge, 3) 3D spatial reasoning, 4) generalization.
@@ 3.1 Environment Interface and Evaluation Protocol
For ease of use and wide compatibility, SmartPlay follows a unified OpenAI Gym interface (Brockman et al., 2016) for all games, with text-based observations, text-based manuals with content as described in Table 1, text-based history describing past actions and observations of length “history length”, and flat categorical actions. Due to the randomness in some games, we recommend running each game multiple times and reporting the average metrics.
Input, manual, action space, rollout length, and trial numbers for each game are specified in Table 1. These settings are fixed and should not be modified. However, future research may require longer history length or more trials for some games. These parameters can be adjusted to suit specific needs, but the changes should be explicitly stated. We provide some recommended values (also used in our experiments) for these parameters in Table 1.
3.2 Example Input
We attach an example input of the game Messenger (Hanjie et al., 2021) as produced by the SmartPlay API. For completeness, we provide example inputs for each game in Appendix B. Note that all directions in SmartPlay are described in “east, south, west, north, above, below”
Instruction Manual:
it loses the game and obtains a reward of -1.
you are being approached by a restricted document that is a robot.
the whale is the main objective.
Observation Example:
You took action Move South.
You (agent) don’t have the message.
You see:
- airplane 7 steps to your south
- fish 13 steps to your south-east
- robot 5 steps to your south-east
In the actual gameplay, SmartPlay API also includes a list of actions for the LLM agent to pick from.
@@ 4.1 Quantitative Analysis
Table 2: Comparison of performance of different LLMs in terms of average score on BanditTwoArmedHighLowFixed-v0, RockPaperScissorBasic-v0, Hanoi3Disk-v0, MessengerL1-v0, MessengerL2-v0, Crafter-v0, MinedojoCreative0-v0. All scores are normalized relative to human performance (unnormalized version in Table 4). GPT-4 variants out-perform other LLMs by significant margins, but still greatly under-perform human baselines. We observe significant performance gaps between SOTA LLMs and human baseline on Hanoi, Crafter, and Minecraft. Hanoi, Crafter challenges planning and reasoning with object dependencies, and Minecraft challenges 3D spatial reasoning.
@@
Open-source LLMs have a long way to go: Open-source LLMs achieves less than half the performance of GPT-4 variants on simple Bandit and Rock-Paper-Scissors tasks, and 1/8 the performance on more challenging tasks. The fine-tuned Vicuna-13b model achieves much worse performance than the base LLAMA-13b.
@@ Appendix A Research Challenges
Reasoning with object dependencies 1. $0\sim 1$-hop dependency 2. $2\sim 3$-hop dependencies 3. multi-hop dependencies 1 1 3 2 3 1
Instruction/Rule Following 1. single game rule 2. <3<3 game rules 3. $5+$ game rules 1 3 2 2 3 2
Planning Ahead 1. <5<5 planning steps 2. $5+$ planning steps 3. concurrent objectives 1 1 3 2 3 1
@@
Randomness 1. no randomness 2. randomness present in game 3. randomness as core mechanism 3 3 1 2 2 2
@@
Error/Mistake Handling 1. not required 2. rollback 3. reasoning and learning 1 1 2 1 3 2
@@
Appendix B Example Inputs
B.1 Bandits
@@
B.2 Rock Paper Scissors
@@ B.6 Minecraft
Appendix C Additional Results

### v4

@@ 1 Introduction
We note $4$ key challenges for intelligent LLM agents not captured in previous benchmarks. First, lots of real-world tasks require an agent to do long-horizon planning. Second, many events are probabilistic and an intelligent agent is expected to understand the odds. Third, an intelligent agent needs spatial reasoning to understand our 3D world. Fourth, when encountered with unseen situations, an intelligent agent should be able to learn from interactions or mistakes.
@@
2 Capabilities Necessary for Intelligent Agents
Borrowing concepts from game design (Koster, 2013), we identify $9$ key abilities important for intelligent LLM agents, and identify multiple degrees for each capability:
  1. a)
    Long text understanding: general LLM capability.
    • We define 4 degrees based on document length and syntactic variations: 1) few fixed lines, 2) few fixed paragraphs, 3) with syntactic variations, 4) and longer than 1 page (500 words).
  2. b)
    Reasoning: multi-hop logical reasoning and deduction, often required for analyzing the interactions of in-game objects or action conditions/dependencies.
    • We define 3 degrees based on reasoning hops: 1) $(0\sim 1)$,  2) $(2\sim 3)$,  3) $(>3)$.
  3. c)
    Instruction/Rule following: follow rules and instructions set by environment or users.
    • We define 3 degrees based on number of game rules: 1) single rule,  2) $(<5)$,  3) $(5+)$
  4. d)
    Planning: long-horizon in-context planning to achieve a complex goal.
    • We define 3 degrees based on planning steps, and concurrent objectives which requires goal prioritization: 1) <5<5 planning steps, 2) $5+$ planning steps, 3) concurrent objectives
  5. e)
    Generalization: Excels at a wide range of tasks.
    • We define 3 degrees based on the variability the game provides: 1) fixed environment, 2) fixed game word with random objectives, 3) procedurally generated game world
  6. f)
    Understanding the odds: analyze and estimate the probability of random events.
    • We define 3 degrees based on the importance randomness in the environment: 1) no randomness, 2) randomness present in game, 3) randomness as core game mechanism
  7. g)
    Learning from interactions: acquire environment knowledge from live interactions.
    • We define 4 degrees based on the number of unique interactions to learn from: 1) no learning required, 2) single interaction, 3) <5<5 interactions, 4) $5+$ interactions
  8. h)
    Error/Mistake handling: recover from mistakes (e.g., correcting from erroneous trajectory).
    • We define 3 degrees based on if mistake handling may be necessary and if additional reasoning and re-planning is necessary: 1) not required, 2) simple rollback corrects error, 3) reasoning and re-planning required to correct error.
  9. i)
    Spatial reasoning: understand our world in 2D/3D. Spatial reasoning is typically required to understand directions and navigate through the game world (e.g., navigating the 2D/3D world).
    1. (a)
      We define 3 degrees based on dimensionality: 1) $0\sim 1$D,  2) 2D,  3) 3D
@@ 3.1 Research Challenges
The SmartPlay benchmark encapsulates a diverse set of challenges that evaluate various AI capabilities, as itemized in Figure 2. For instance, Bandits primarily focuses on understanding the odds, requiring minimum text understanding and rule-following. On the other hand, Rock Paper Scissors uniquely puts an emphasis on understanding the odds and multiple game rules. Hanoi presents an advanced setting for object dependency reasoning, strategic planning, and handling mistakes. Messenger puts challenge on 2D spatial reasoning, reading syntactic variations and conducting multi-hop reasoning. Meanwhile, Minecraft offers a unique challenge in 3D spatial reasoning and generalization within a procedurally generated world. We hope the SmartPlay benchmark would serve as a tool for identifying these nuanced gaps and directing future research.
@@
Multi-hop Reasoning 1. $0\sim 1$-hop 2. $2\sim 3$-hop 3. multi-hop
Instruction/Rule Following 1. single game rule 2. <5<5 game rules 3. $5+$ game rules
Planning 1. <5<5 planning steps 2. $5+$ planning steps 3. concurrent objectives
@@
Understanding the Odds 1. no randomness 2. randomness present in game 3. randomness as core mechanism
@@
Error/Mistake Handling 1. not required 2. rollback only 3. reason and replan
@@
Figure 2: We identify a set of 9 important capabilities (section 2) for an intelligent agent. We identify different degrees of challenge for each capability as shown on the left. Each game in SmartPlay challenges unique set of capabilities at different degrees, as shown in the spider charts. We include numerical values of the spider plots in Table 3.
@@ 3.2 Two Armed Bandits
The LLM agent is provided two slot machines with hidden pre-defined reward probabilities $p_{1},p_{2}$. For slot machine $i$, the reward for the two possible out-comes are: $r_{i}$ for pay-off event and $r_{i}$ for no-pay-off event. The goal of the game is to find the arm with better return and maximize the reward over the course of 50 rounds. The human written manual informs the LLM of the number of slot machines (two) and the objective.
@@ 3.3 Rock Paper Scissors
Same rules as the famous zero-sum game Rock Paper Scissors222wikipedia.org/wiki/Rock_paper_scissors.
The LLM agent plays against a hand-coded opponent that follows a hidden pre-defined strategy with probabilities $p_{1},p_{2},p_{3}$ for rock, paper, and scissors respectively. The scores for winning under each action is pre-defined and revealed to the LLM as $s_{1},s_{2},s_{3}$. The human written manual provides instruction on the possible actions and how the win/draw/lose of each round is calculated.
An agent must keep track of win/losses from its past roll-outs to analyze the opponent behavior, and then exploit the opponent to maximize payoff. Overall, the challenges include: 1) long text understanding, 2) understanding the odds, 3) learning from interactions, 4) instruction following.
@@ 3.4 Tower of Hanoi
The Tower of Hanoi requires the agent to think strategically and plan ahead, and put strict requirements on the LLM agents’ ability to understand and follow the rules of the game. The game can become more challenging if the agent makes a mistake. Sometimes, an agent may have to undo several moves to correct an error. Overall, the challenges include: 1) planing, 2) reasoning, 3) error handling.
@@ 3.5 MESSENGER
To succeed in MESSENGER, an agent must first relate entities and dynamics of the environment to their reference synonyms in the manual, identify message and goal objects, and navigate to bring the message to the goal while avoiding the enemy. The manual, by design, is challenging to understand even for human readers. Level 1 primarily challenges the agent’s 1) long text understanding and 2) generalization . Level 2 includes additional challenge on the agent’s 3) reasoning, and 4) 2D spatial reasoning. Level 3 increases difficulty by adding distraction objects.
The original manuals provided by Hanjie et al. (2021) contain descriptions of the entities and world dynamics obtained through crowd-sourced human writers. We augment the manual with a specification on the game objective, and an “advice” for LLM agent to first identify goal objects and then approach its objective. The “advice” reduces the difficulty of the hard-to-parse manual.
@@ 3.6 Crafter
We provide the “context” string from Wu et al. (2023c) as the manual, generated by parsing the source-code of (Hafner, 2021). The “context” string has been shown to greatly improve performance of GPT-4 and text-davinci-003 on Crafter (Wu et al., 2023c).
To succeed in Crafter an LLM agent has to first understand and master a variety of reusable skills composed of 17 actions. The agent needs to learn to navigate through up to 5 2D terrains (biomes), avoiding obstacles and dangerous creatures. The agent also needs to collect different resources and craft more advanced weapons/tools to unlock more skills and achievements, while at the same time balance crafting goals with survival goals like maintaining health, thirst, food, and rest (Hafner, 2021). Overall, the challenges include: 1) 2D spatial reasoning, 2) mistake handling, 3) long text understanding, 4) planning, 5) generalization, 6) correcting from mistakes. Interestingly, the “context” string does not capture all information necessary to succeed in the game, i.e., it requires 2 woods to craft the crafting table, and 8 stones to craft the furnace. The agent has to 7) learn from interaction.
@@ 3.7 Minecraft
To succeed in the creative “find” tasks, a LLM agent has to have enough domain knowledge about different biomes in Minecraft, and be able to correlate visual observation (text description of visual world) with domain knowledge, and navigate in a 3D environment. Overall, the challenges include: 1) planning, 2) domain knowledge, 3) 3D spatial reasoning, 4) generalization.
@@ 4.1 Environment Interface and Evaluation Protocol
For ease of use and wide compatibility, SmartPlay follows a unified OpenAI Gym interface (Brockman et al., 2016) for all games, with text-based observations, text-based manuals with content as described in Table 1, text describing historical actions and observations covering past steps of length “history length”, and flat categorical actions. Due to the randomness in some games, we recommend running each game multiple times and reporting the average metrics.
Input, manual, action space, rollout length (the maximum environment steps allowed by each game), and trial numbers for each game are specified in Table 1. These settings are fixed and should not be modified. However, future research may require longer history length or more trials for some games. These parameters can be adjusted to suit specific needs, but the changes should be explicitly stated. We provide recommended values (also used in our experiments) for the parameters in Table 1.
For completeness, we provide example inputs for each game in Appendix C. Note that all directions in SmartPlay are described in “east, south, west, north, above, below” In the actual gameplay, SmartPlay API also includes a list of actions for the LLM agent to pick from.
@@ 5.1 Quantitative Analysis
Table 2: Comparison of performance of different LLMs in terms of average score on BanditTwoArmedHighLowFixed-v0, RockPaperScissorBasic-v0, Hanoi3Disk-v0, MessengerL1-v0, MessengerL2-v0, Crafter-v0, MinedojoCreative0-v0. All scores are normalized relative to human performance (unnormalized version in Table 4). GPT-4 variants out-perform other LLMs by significant margins, but still greatly under-perform human baselines. We observe significant performance gaps between SOTA LLMs and human baseline on Hanoi, Crafter, and Minecraft. Hanoi, Crafter challenges planning and reasoning, and Minecraft challenges 3D spatial reasoning.
@@
Open-source LLMs have a long way to go: Open-source LLMs achieves less than half the performance of GPT-4 variants on simple Bandit and Rock-Paper-Scissors tasks, and 1/8 the performance on more challenging tasks. The fine-tuned Vicuna-13b model performs much worse than the base LLAMA-13b.
@@ Appendix A Research Challenges
Reasoning 1. $0\sim 1$-hop 2. $2\sim 3$-hop 3. multi-hop 1 1 3 2 3 1
Instruction/Rule Following 1. single game rule 2. <5<5 game rules 3. $5+$ game rules 1 3 2 2 3 2
Planning 1. <5<5 planning steps 2. $5+$ planning steps 3. concurrent objectives 1 1 3 2 3 1
@@
Understanding the Odds 1. no randomness 2. randomness present in game 3. randomness as core mechanism 3 3 1 2 2 2
@@
Error/Mistake Handling 1. not required 2. rollback only 3. reason and re-plan 1 1 2 1 3 2
@@
Appendix B Minecraft Visual Descriptor
Figure 4: Left: Raw Minecraft environment observation Right: Segmentation map of the environment observation. as detected by lidar rays in MineDojo (Fan et al., 2022).
The raw ground truth MineDojo (Fan et al., 2022) is a block level matrix (2D matrix for lidar rays and 3D matrix for surrounding blocks), which is very hard for human or LLMs to comprehend. Inspired by Wu et al. (2023c), we adopt a directional visual description scheme to encode the scene observation in text. Specifically, we first run connected component algorithm to group the same blocks that are connect into groups, and then describe the group’s relative position to the agent based on the closest block from the group. For example, an observation in Figure 4 will be described as:
You see:
 - grass, 3.87 blocks away to your bottom, taking 38% of screen
 - plant, 22.10 blocks away to your bottom-right, taking 2% of screen
 - water, 33.88 blocks away to your bottom-left, taking 1% of screen
 - wood, 18.76 blocks away to your bottom-left, taking 0% of screen
 - flower, 32.98 blocks away to your right, taking 0% of screen
 - wood, 24.50 blocks away to your left, taking 0% of screen
 - leaves, 11.63 blocks away to your left, taking 6% of screen
 - leaves, 56.55 blocks away to your right, taking 1% of screen
Appendix C Example Inputs
C.1 Bandits
@@
C.2 Rock Paper Scissors
@@ C.6 Minecraft
Appendix D Additional Experimental Results
D.1 Human Baseline
3 players (including the authors) who are very familiar with the environments and API played the games through the SmartPlay interface. Each human player performed 3 rounds of Bandit, RPS; 1 round of Hanoi, Crafter, Minecraft; 5 rounds of MessengerL1, MessengerL2. We report the final average score over all trials and all players.
D.2 Normalized Human Score
Given the game score of an LLM on game $g$, $s_{g}^{\text{(raw)}}$, we compute normalized human score $s_{g}$ from the human baseline on $g$, $s_{g}^{\text{(human)}}$, and the minimum possible game score $s_{g}^{\text{(min)}}$:
$s_{g}=\frac{s_{g}^{\text{(human)}}-s_{g}^{\text{(raw)}}}{s_{g}^{\text{(human)}}-s_{g}^{\text{(min)}}}$
D.3 Raw scores